OpenStudy (anonymous):

What is the slope of the line between (-4, 4) and (-1, -2) how do i do problems like these not really looking for a answer i just want to learn it

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-4\quad ,&4)\quad &(-1\quad ,&0-2) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

\[slope = \frac{ (y _{2} -y _{1})}{ (x _{2}- x _{1}) }\]

OpenStudy (jdoe0001):

hmm \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-4\quad ,&4)\quad &(-1\quad ,&-2) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)